OpenStudy (anonymous):

The sum of two numbers is 20 and their difference is 2, write system of linear equations and solve the?

OpenStudy (anonymous):

10 and 4 sorry

OpenStudy (anonymous):

x + y = 20 x - y = 2 -----------add 2x = 22 x = 11 x + y = 20 11 + y = 20 y = 20 - 11 y = 9 check... x + y = 20 11 + 9 = 20 20 = 20 (correct) check... x - y = 2 11 - 9 = 2 2 = 2 (correct) any questions ?

OpenStudy (anonymous):

wait...what is the 10 and 4 ? Is it supposed to be x + y = 10 and x - y = 4 ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x + y = 10 x - y = 4 -----------add 2x = 14 x = 7 x + y = 10 7 + y = 10 y = 10 - 7 y = 3 check... x + y = 10 7 + 3 = 10 10 = 10 (correct) check.. x - y = 4 7 - 3 = 4 4 = 4 (correct)

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

how would i graph it

OpenStudy (anonymous):

I am sorry...I am not good at graphing. I do know that to find the x intercept of a line is to sub in 0 for y and to find the y intercept, sub in 0 for x.